David has a lot of homework to do. He starts his reading homework at 3:45 and ends at 4:30. Then he does math from 4:30 until 5:00. Lastly, he studies for a science test from 5:00-5:30. How much total time did David spend on his homework and studying?
Answer by Cromlix(4381)

Hi there,
Reading = 4.30 - 3.45 = 45 mins
Maths = 5.00 - 4.30 = 30 mins
Science study = 5.30 - 5.00 = 30 mins
Total of study + homework = 45 + 30 + 30 = 1 hour 45 mins
David spent 1 hour 45 minutes doing homework and studying.
Hope this helps :-)